OpenStudy (anonymous):

PLEASE HELP Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now? Numerical Answers Expected!

OpenStudy (lgbasallote):

let tim's present age = x coach = 10 + x (10 + x) - 3 = 2(x-3) 7 + x = 2x -6 x = 13 10 + x = 23